Jim and Juan are going to drive their motorcycles from Gastonia to Charleston. Jim leaves at 8:00 am but Juan can not leave until 8:30 am. If Jim drives at 50 mph and Juan drives at 55 mph, at what time will Juan catch up with Jim?
Answer by checkley77(12844)

50T=55(T-.5)
50T=55T-27.5
50T-55T=-27.5
-5T=-27.5
T=-27.5/-5
T=5.5 HOURS AFTER 8:00 AM THEY WILL MEET OR @ 1:30 PM.